### 30 Mar 2026 v1

# Gravitational Redshift of Light and the Heisenberg Uncertainty Principle

Abstract
Empirical observations together with theoretical analyses are being used to argue that the classical phenomenon of gravitational redshift—namely, the redshift of light in a static gravitational potential—may be in tension with the Heisenberg uncertainty principle. In particular, in the Pound–Rebka experiment, the emitter–absorber pair interact with each other by exchanging a Mössbauer photon, a process that is fully describable within quantum mechanics but how the photon interacts with the spacetime metric remains unclear. Since the uncertainty principle is incompatible with local realism, as per general relativity, we propose to study continuous-variable photonic entanglement within the Einstein–Podolsky–Rosen (EPR) framework in a weak gravitational field. We outline a thought experiment, realizable on the surface of the Earth, that could shed further light on this problematic seamline between general relativity and quantum mechanics in the weak-field regime.
Keywords: Gravitational redshift of light, Heisenberg uncertainty, photonic entanglement and EPR
I Introduction
Gravitational redshift (GRS) can be derived from first principles and was predicted by Einstein, on the basis of the equivalence principle, prior to the formulation of general relativity (GR) [1]. Subsequent analyses have argued that, unless the energy of a photon traversing a static gravitational potential decreases, the first law of thermodynamics would be violated [^1] The argument proceeds as follows. Consider an excited atom at point 2 with energy, $E_{2}=m_{A}c^{2}+\epsilon=m_{A}^{*}c^{2}$, fall in a gravitational field to a point 1. Gravity does work, $m_{A}^{*}\Delta\Phi$. Let the atom radiate a photon at point 1, $\epsilon=h\nu_{1}$, thereby reducing the atom’s energy to $m_{A}c^{2}$. Now the atom is taken back to point 2. Gravity does work, $m_{A}\Delta\Phi$. If the emitted photon’s frequency is not changed it could excite an identical atom causing a cyclic process in which work is done and energy is produced, $\oint dW=\left(m_{A}^{*}-m_{A}\right)\Delta\Phi=\epsilon\frac{\Delta\Phi}{c^{2}}$. Unless the energy of the photon is decreased by GRS such that $h(\nu_{1}-\nu_{2})=\epsilon\frac{\Delta\Phi}{c^{2}}$, the first law of thermodynamics would be violated [2]. [2]. As a quantum object, any change in the photon’s energy and momentum must comply with the Heisenberg uncertainty principle (HUP). However, cross-verification of the momentum–position uncertainty relation against empirical data reveals a persistent tension. Reformulating this relation solely in terms of wavelength leads to paradoxical constraints that conflict with GR. Furthermore, because photon momentum is not conserved, we develop a model from first-principles that independently substantiates these inconsistencies. Taken together, these results indicate an incompatibility between GR and QM that manifests itself even within the weak-field regime.
II Tension between General Relativity and Quantum Mechanics
In GR, time dilation causes a clock based on an atomic or molecular transition at ground level to tick differently than one held aloft at height $z$ above the ground. Accordingly, the frequency of a photon at altitude $z$ is related to that at ground level by:
$\Delta\nu_{z}=\nu_{z}-\nu_{0}=\frac{\nu_{0}gz}{c^{2}\left(1+\frac{z}{R_{E}}\right)}\approx\nu_{0}z\times\left(1.09\times 10^{-16}\right).$ (1)
Where $R_{E}$ is Earth’s radius and $z$ is the altitude measured in meters [3]. Equation (1) can be generalized to describe the frequency shift between any two altitudes $z$ and $z^{\prime}$, on Earth’s surface provided that $z^{\prime},z\ll R_{E}$,
$\Delta\nu_{z^{\prime}z}=\left(\nu_{z^{\prime}}-\nu_{z}\right)\approx\nu_{0}(z^{\prime}-z)\times 1.09\times 10^{-16}.$ (2)
II.1 GRS and the Momentum–Position uncertainty
In QM, the momentum of a photon is given by:
$p=\frac{h}{\lambda},$
so the change in wavelength implied by Equation (1) corresponds directly to a change in the photon’s momentum. Moreover, the expectation values of the photon’s momentum and position become interdependent and are constrained by the HUP. For a photon propagating in the $z$-direction above ground, the uncertainties in position and momentum satisfy:
$\Delta p_{z}\Delta z\geq\frac{\hbar}{2},$
where $\Delta z$ denotes the uncertainty in the photon’s position along the propagation axis and $\Delta p_{z}$ the corresponding uncertainty in its momentum.
We can recast these uncertainties in terms of the photon wavelength as following:
$\Delta p$ $=\sqrt{\left(\delta\lambda\times\frac{dp}{d\lambda}\right)^{2}}=h\frac{\delta\lambda}{\lambda^{2}},$ (3a)
$\Delta z$ $=\kappa\lambda.$ (3b)
where $\kappa>0$ is a dimensionless parameter that may depend on the resolution, $\left(\delta\lambda/\lambda\right)$, of the wavelength (e.g. the natural bandwidth at FWHM). Categorically, any characteristic length scale can be expressed in units of the photon’s wavelength. Examples include the coherence length $L_{c}\approx\lambda/(\delta\lambda/\lambda)$, where $\delta\lambda/\lambda$ denotes the spectral resolution; and the beam waist $w_{0}=\lambda/(\pi\theta)$, with $\theta$ the beam’s divergence angle. If the photon is modeled as a wavepacket of spatial width $\Delta z=\sigma$, and thereby a minimum-uncertainty state, a monochromatic wavepacket cannot exhibit a spread smaller than its wavelength $\lambda$. Consequently, expressing the positional uncertainty as $\Delta z=\kappa\lambda$ requires $\kappa\geq 1$. Generally, the positional uncertainty may be expressed as a function of the spectral resolution $\delta\lambda/\lambda$ and the photon’s wavelength:
$\Delta z=\lambda f\left(\frac{\delta\lambda}{\lambda}\right).$
Expanding $f$ in a Taylor series about $\delta\lambda/\lambda\to 0$, we obtain:
$\Delta z=\lambda\left[f(0)+f^{\prime}(0)\left(\frac{\delta\lambda}{\lambda}\right)+f^{\prime\prime}(0)\left(\frac{\delta\lambda}{\lambda}\right)^{2}+\ldots\right].$
To leading order this reduces to Equation (3b)222:Consider the following, we can recast the position in terms of the wavelength as, $z=q\lambda$, where $q$ is the number of wavelengths that fit into the distance $z$. For one wavelength we have, $\delta z_{1}\sim\delta\lambda\equiv(\delta\lambda/\lambda)\lambda$, with $\delta\lambda/\lambda$, the resolution of the wavelength, usually related to the FWHM. The RMS of these $q$ deviations is, $\Delta z_{\mathrm{rms}}=\sqrt{\sum\frac{1}{q}\left(\frac{\delta\lambda_{i}}{\lambda}\right)^{2}\lambda^{2}}=\left(\frac{\delta\lambda}{\lambda}\right)\lambda$, with a maximum uncertainty, $\Delta z_{\max}=\sqrt{q\left(\frac{\delta\lambda}{\lambda}\right)^{2}\lambda^{2}}=\sqrt{q}\left(\frac{\delta\lambda}{\lambda}\right)\lambda$. The resolution is a numerical factor and the numerical value is renamed as, $\sqrt{q}\left(\frac{\delta\lambda}{\lambda}\right)\equiv\kappa$, leading to $\Delta z=\kappa\lambda$.
$\Delta z\simeq\lambda f(0)=\kappa\lambda.$
Combining (3a) & (3b) we end up with:
$\eta^{\mathrm{HUP}}$ $=\Delta p\times\Delta z=h\kappa(\delta\lambda/\lambda)\geq\frac{h}{2\pi},$ (4a)
$\frac{1}{2\kappa\pi}$ $\leq(\delta\lambda/\lambda).$ (4b)
Note that the Planck constant, $h$, drops out of Equations (4), underscoring their intrinsically classical character. The wavelength shift of a photon ascending from Earth’s surface to an height $z$ is calculated in GR with the Schwarzschild metric [5]:
$\frac{\Delta\lambda}{\lambda}\approx\frac{2M_{E}G}{c^{2}R_{E}^{2}}=\frac{gz}{c^{2}}.$ (5)
Also note that Equation (5a) can be reformulated as:
$h\frac{\Delta\lambda}{\lambda}=h\frac{gz}{c^{2}}.$ (6)
Evidently, Equations (5) presuppose a well-defined value for the photon’s momentum $h/\lambda$ along with a definite position $z$. However, the HUP forbids the simultaneous assignment of sharp values to these conjugate variables. We will demonstrate that the empirical data from the Pound and Rebka experiment [3] are in tension with Equations (4). The Pound–Rebka experiment employed a Mössbauer gamma-ray source with an energy of $14.4\,\mathrm{keV}$, corresponding to a wavelength of approximately $\lambda\approx 8.63\times 10^{-11}\,\mathrm{m}$. The full width at half maximum (FWHM) of the spectral line was about $4.7\times 10^{-9}\,\mathrm{eV}$ [3]. This yields a spectral resolution of $(\delta\lambda/\lambda)\approx 0.32\times 10^{-12}$, whereas the gravitational redshift (GRS) measured in the experiment was $\Delta\lambda/\lambda\approx 2.5\times 10^{-15}$. For Equation (4b) to hold true we would need a value of:
$\kappa\geq(2\pi\delta\lambda/\lambda)^{-1}\approx 2.5\times 10^{11}:,$ (7)
which deems the estimated uncertainty in position for a wavelength of $\lambda\approx 8.63\times 10^{-11}\,\mathrm{m}$ to be:
$\Delta z=\kappa\lambda\geq 21.5\,\mathrm{m}.$ (8)
This value is comparable to the height of the tower ($z=22.5\,\mathrm{m}$) in the Jefferson Laboratory at Harvard University, where the experiment was conducted. This does not contradict the coherence length of a Mössbauer photon, $L_{c}\approx\lambda/(\delta\lambda/\lambda)\sim 260\,\mathrm{m}$, which simply reflects the probability of finding the photon over that distance. The problem, rather, is how to conceptually reconcile the uncertainty—or, for that matter, the coherence length—with the strong height dependence implied by Equation (1).
Quantum theory describes the Mössbauer emission mechanism as a single quanta:
$|1\rangle=\int d\omega\,f(\omega)\,|1_{\omega}\rangle,$
with a Lorentzian distribution around $\omega_{0}$, and a width $\Gamma$,
$f(\omega)\propto[(\omega-\omega_{0})+i\Gamma/2]^{-1}:$
As the linewidth is extremely narrow, on the order of $10^{-9}\,\text{eV}$, the emitted pulse has a correspondingly long lifetime of about $141\,\text{ns}$. According to the Weisskopf–Wigner theory, the photon inherits the exponential decay profile of the nuclear transition, leading to a time-domain wavefunction of the form [4]:
$\psi(t)\propto e^{-i\omega_{0}t}\,e^{-t/(2\tau)}\,\Theta(t).$
Here, $\Theta(t)$ represents the sharp onset of the emission, while the decay is purely exponential. As a result, the photon is emitted over a characteristic time $\tau$, so its wavepacket length is $L\sim c\tau\approx 40\,\mathrm{m}$. This is smaller than the coherence length $L_{c}$ by a factor of $2\pi$, reflecting the ratio $h/\hbar$. Obviously, this description is not local—neither in space nor in time—since the leading edge of the wavepacket experiences substantially more time dilation than its tail. It may be tempting to imagine that a wavepacket “samples” multiple spacetime regions, each with its own redshift, or that its leading and trailing components persist in quantum superposition. But within a QFT framework, such a picture is untenable. Any putative component of the wavepacket that experiences a distinct gravitational redshift must interact locally with the corresponding region of the background geometry. This immediately elevates those supposed “components” to the status of independent physical degrees of freedom, not mere substructures of a single coherent excitation. The unavoidable consequence is that gravitational time dilation forces different segments of the wavepacket to accumulate inequivalent phases, and this phase dispersion is flatly incompatible with the preservation of coherence.
One may reasonably argue that the measurement protocol in the Pound–Rebka experiment [3] is correctly described by an absorption-probability functional that does not violate the Heisenberg uncertainty principle. For example, consider the description of the fractional counting rate in the experiment as a Lorentzian probability distribution, where the linewidth $\Gamma$ enters as the width of the Lorentzian while the gravitational redshift produces a deterministic shift of its center, and the Doppler-compensating velocity, $v$, is used to restore resonance:
$R(v)\propto\frac{\Gamma^{2}}{(h\nu\,v/c-\Delta E_{g})^{2}+\Gamma^{2}}.$
For this to hold, the absorber must resonate with the emitted photon, such that the absorber’s probability effectively selects the portion of the photon’s spectrum that overlaps with its resonance:
$P_{\mathrm{abs}}\propto\left|\int d\omega\,A_{\mathrm{em}}(\omega)A_{\mathrm{abs}}^{*}(\omega)\right|^{2}.$
According to quantum mechanics, the emitter–absorber pair interacts by exchanging a Mössbauer photon, yet how this photon couples to the spacetime metric—specifically in a way that leads to an energy loss—remains obscure. In the quantum picture, the photon mediates a coherent, recoil-free exchange with spectral properties fixed by the transition dynamics, whereas general relativity treats it as a classical test particle whose frequency is deterministically shifted by the gravitational potential via time dilation. Reconciling these descriptions is nontrivial. If the photon’s energy–momentum is not well defined during flight but becomes definite only upon measurement, the cumulative redshift predicted by GR becomes conceptually puzzling. Conversely, if the gravitational field acts continuously on the photon, this presupposes a quasi-classical trajectory with well-defined dynamical variables, in tension with the Heisenberg uncertainty principle. This reveals a deeper gap in our understanding of how quantum systems couple to classical spacetime, even in weak fields.
The observed agreement with the GR prediction, $\Delta\nu_{\mathrm{obs}}/\Delta\nu_{\mathrm{GR}}\simeq 1$, only sharpens the issue: according to Equation (1), each additional meter of elevation should degrade this accuracy dramatically. What, then, is the operational meaning of the uncertainty principle for a photon wavepacket propagating in Earth’s gravitational field, and how does this comport with the deterministic structure of general relativity? If the photon’s momentum is not a real, well-defined property across its entire wavepacket during its ascent, how can the cumulative effect implied by Equation (5) arise? In this view, spacetime effectively acts as a continuous probe of the photon’s spatial and temporal properties—an assumption that implicitly reinstates a form of local realism. The fact that the measurement procedure respects the uncertainty principle does not resolve the paradox: gravitational redshift appears to rely on a quasi-classical, locally realistic account of the photon’s dynamical variables, in direct tension with the quantum-mechanical prohibition against their simultaneous definiteness.
II.2 GR Imposed Constraints on $\kappa$
The gravitational redshift of a photon emitted from the surface of a massive body at radius $r_{E}$ and observed at earth at a distance $r_{O}\gg r_{E}$, is calculated by the Schwarzschild metric as:
$\frac{\lambda_{O}}{\lambda_{E}}=\frac{\sqrt{1-\frac{2GM}{r_{O}c^{2}}}}{\sqrt{1-\frac{2GM}{r_{E}c^{2}}}}\underset{r_{0}\to\infty}{\cong}\frac{1}{\sqrt{1-\frac{2GM}{r_{E}c^{2}}}}.$ (9)
The quantity $\epsilon=\frac{2GM}{r_{E}c^{2}}$ is the ratio between the Schwarzschild radius, $\frac{2GM}{c^{2}}$, and the physical radius of the star, $r_{E}$. The fractional shift is therefore:
$\frac{\left(\lambda_{O}-\lambda_{E}\right)}{\lambda_{E}}=\frac{\Delta\lambda}{\lambda}=\frac{1}{\sqrt{1-\epsilon}}-1.$ (10)
GRS values increase as $\epsilon$ approaches the size of its black hole radius. Indeed as $\epsilon\rightarrow 1$, the shift becomes infinite. An infinite shift is consistent with the fact that light cannot escape the horizon of a black hole.
In the following, we shall assume that the numerical values predicted by Equation (9), as they may be realized through a quantum measurement protocol, are typically smaller than the intrinsic resolution set by the linewidth $\Gamma$. That is, for a measurement procedure involving large statistics, gravity shifts the center of the distribution, whereas the linewidth reflects the underlying quantum uncertainty. However, for the interaction of a single photon with spacetime, the gravitationally induced shift would be obscured by this intrinsic uncertainty. Under the assumption that gravity acts on individual photons according to GR, we express the relation between gravitational redshift and the Heisenberg uncertainty principle as:
$\frac{1}{2\kappa\pi}\leq\frac{\Delta\lambda}{\lambda}=\frac{1}{\sqrt{1-\epsilon}}-1.$ (11)
which provides a lower bound on the fractional wavelength shift due to the HUP and an upper bound due to GR. For large $\epsilon$ values, the wave shifts are in accord with HUP as no special restriction is necessary for $\kappa$ in Equation (10). However, for smaller values of $\epsilon$, $\kappa$ cannot be a mere a constant but must be a monotonously decreasing function of $\epsilon$:
$\kappa(\epsilon)\geq\frac{\sqrt{1-\epsilon}}{4\pi(1-\sqrt{1-\epsilon})}.$ (12)
The postulated ad hoc function, $\kappa(\epsilon)$, illustrated in Fig. 1, is neither a component of general relativity nor of quantum mechanics. It must be dimensionless and simultaneously intrinsic to the structure of spacetime but cannot be measured. We argue that this suggests incompatibility, in spacetime at large, between GR and QM.
Figure 1: GRS as calculated from Equation 9. The insert is a log-log plot of Equation 11.
II.3 A Toy Model for Photon Propagation Under GRS
As indicated by Equation (2) the energy and therefore the momentum of a photon traveling in a gravitational field is not conserved, as $\delta E_{zz^{\prime}}=h\Delta\nu_{z^{\prime}z}$. Indeed, QM states that the momentum must be proportional to the gravitational potential:
$\dot{p}=[p,H]=[p,\Phi]=-i\hbar\frac{\partial\Phi}{\partial z}\neq 0.$ (13)
This holds true in both the Schrödinger and Heisenberg pictures, where Equation (12) applies to the observable eigenvalue of the momentum operator, $p$, per Ehrenfest’s theorem:
$\partial_{t}\langle p_{z}\rangle=-\langle\partial_{z}\Phi\rangle.$
This allows the construction of a model describing the motion of a photon in a static gravitational field as a succession of local interactions each of which resulting in an infinitesimal local shift of its wavelength.
For a photon traveling the distance $\Delta z=z^{\prime}-z$, in a static gravitational field, the change in momentum from $p_{z}=h/\lambda_{z}$, to, $p_{z^{\prime}}=h/\lambda_{z^{\prime}}=h/(\lambda_{z}+\Delta\lambda_{zz^{\prime}})$, at $z^{\prime}$, is:
$\Delta p_{zz^{\prime}}=h\left(\frac{1}{\lambda_{z^{\prime}}}-\frac{1}{\lambda_{z}}\right)=-h\frac{\Delta\lambda_{zz^{\prime}}}{\lambda_{z}(\lambda_{z}+\Delta\lambda_{zz^{\prime}})}.$ (14)
Dividing both the numerator and denominator by $\lambda_{z}$ we obtain:
$\Delta p_{zz^{\prime}}=\left(\frac{-h}{\lambda_{z}}\right)\left(\frac{\Delta\lambda_{zz^{\prime}}/\lambda_{z}}{1+\Delta\lambda_{zz^{\prime}}/\lambda_{z}}\right).$ (15)
For $\Delta\lambda_{zz^{\prime}}/\lambda_{z}\ll 1$, and in first order, we have:
$\Delta p_{zz^{\prime}}\cong\left(\frac{-h}{\lambda_{z}}\right)\left(\frac{\Delta\lambda_{zz^{\prime}}}{\lambda_{z}}\right)=\frac{-h\Delta\lambda_{zz^{\prime}}}{\lambda_{z}^{2}}.$ (16)
The momentum-position uncertainty relation follows Equations (3) but at the same time cannot be larger than $\Delta p_{zz^{\prime}}\times(z^{\prime}-z)$. Here again we cast the distance traveled in term of the wavelength as $\Delta z=z^{\prime}-z=\varkappa\lambda_{z}$, to arrive at the following inequality:
$\frac{1}{2\pi}\leq\frac{\varkappa\Delta\lambda_{zz^{\prime}}}{\lambda_{z}}.$ (17)
This inequality is strongly dependent on the constant, $\varkappa$, but since the relative wavelength shift is very small, $\Delta\lambda_{zz^{\prime}}/\lambda_{z}\approx 10^{-15}$, for Equation (16) to hold true, $\varkappa$ has to be of the order of $\varkappa\sim O(10^{15})$, which is highly unlikely. This result sharpens the paradox suggested above (Section II.A). At each position during its propagation through a static gravitational field, while losing energy via GRS, the photon has to have real simultaneous continuous variables, momentum and position, while at the same time it must succumb to the uncertainty principle which states the opposite.
III HUP and Entanglement
In the sections above, evidence for tension between GRS and HUP was shown both from the empirical and theoretical aspects. This, albeit conceptual tension, suggests that already at the weak gravitational field, gravitational redshift may be incompatible with the Heisenberg uncertainty principle. If the Heisenberg assertion that momentum and position cannot assume simultaneously physical realities is false, then the picture painted by general relativity would explain GRS, even for a quantum object such as a photon, rather classically. The very fact that the uncertainty relations as per Equations (4,5), is shedding the Planck’s constant, $h$, encapsulates succinctly this message. Namely, the uncertainty relations lose their quantum context and wholly conform to a classical picture.
Photonic entanglement deems quantum mechanics a nonlocal theory, in contradiction with general relativity and the equivalence principle. This is best described by the EPR paradox [6]. Entangled photons that undergo gravitational redshift would entangle spacetime itself, thereby turning it into a quantum entity. So far, as has been observed in numerous tests of general relativity, spacetime adheres to its local character. Extending the argument of local realism from single quanta photons to entangled photons would mean that the entangled pair violates the uncertainty principle with respect to the independent reality of their positions and momenta in spacetime. No “steering” [^3] Schrödinger introduced the concept of entanglement or “steering” to express the nonlocal effect of part of a system on another part of it, in what seems to be an action at a distance. information has to travel through spacetime faster than the speed of light in order to maintain the correlation between the members of the pair. The initial correlation would be altered by the local interaction of each photon of the pair with spacetime. As no measurement is performed the “collapse of the wavefunction” is avoided altogether.
At this point we note that the observed entanglement in the polarization of photons does not cause any tension between quantum mechanics and general relativity, as no spin-dependent effects are expected in a weak gravitational field. This is because intrinsic spin is equivalent to ordinary angular momentum. If this were not the case, the equivalence principle would be violated, however, thus far all empirical evidence support it [7]. The entanglement in helicity states is therefore a feature of quantum mechanics alone. One can invoke the second law of thermodynamics in providing contextual arguments for why quantum entanglement in helicity states does not pose a fundamental difficulty for general relativity while GRS does. Since the entangled pair remains a closed system with respect to their coupled helicity states, and no energy is dispersed through spin–gravity interaction, the system’s entropic information remains unchanged [8]. This, however, does not hold for the photons’ momenta, as indicated in Equation (12). The energy dissipated when a photon undergoes gravitational redshift necessarily leads to an increase in entropy.
Assume a pair of entangled photons, one traveling a vertical distance $z$ above Earth’s surface and undergoing gravitational redshift. Its entangled partner remains at the surface, ($z=0$), so its wavelength does not shift. How does GRS affect the entangled partner that doesn’t experience it directly?
To this end we propose to test the effect with a thought experiment utilizing a set of EPR entangled beams on Earth’s surface. Entanglement-based EPR arguments presuppose a single global time parameter $t$ with respect to which joint observables such as:
$\widehat{X}_{-}=\widehat{x}_{A}(t)-\widehat{x}_{B}(t)$
and
$\widehat{P}_{+}=\widehat{p}_{A}(t)+\widehat{p}_{B}(t)$
are defined as simultaneous measurements on spatially separated subsystems. In a curved spacetime, by contrast, each subsystem follows a worldline, $x_{j}^{\mu}(\tau_{j})$, with its own proper time, $\tau_{j}$, related to any chosen coordinate time $t$ by:
$d\tau_{j}=\sqrt{-g_{\mu\nu}(x_{j})\dot{x}_{j}^{\mu}\dot{x}_{j}^{\nu}}\,dt,$
which in the weak-field limit reduces to:
$d\tau_{j}\simeq\left(1+\Phi_{j}/c^{2}\right)dt.$
The local mode operators evolve as:
$\widehat{a}_{j}(\tau_{j})=\widehat{a}_{j}(0)e^{-i\omega_{j}\tau_{j}},\qquad\widehat{a}^{\dagger}_{j}(\tau_{j})=\widehat{a}^{\dagger}_{j}(0)e^{-i\omega_{j}\tau_{j}},$
so that expressing both subsystems in a single time parameter $t$ induces unequal effective frequencies:
$\omega_{j}^{\mathrm{eff}}=\omega_{j}\left(1+\Phi_{j}/c^{2}\right),$
and, at the level of the canonical quadratures [9]:
$\widehat{x}_{j}=\frac{1}{\sqrt{2}}\left(\widehat{a}_{j}+\widehat{a}_{j}^{\dagger}\right),\qquad\widehat{p}_{j}=\frac{1}{\sqrt{2}}\left(\widehat{a}_{j}-\widehat{a}_{j}^{\dagger}\right),$
a nontrivial symplectic map:
$(\widehat{x}_{j},\widehat{p}_{j})\mapsto(\sqrt{\zeta_{j}}\,\widehat{x}_{j},\,\widehat{p}_{j}/\sqrt{\zeta_{j}})$
with $\zeta_{j}\simeq\left(1+\Phi_{j}/c^{2}\right)$.
The global EPR observables $\widehat{X}_{-}$ and $\widehat{P}_{+}$ become explicitly gravity-dependent constructs whose canonical form is not invariant under changes of gravitational potential, even though the underlying bipartite state should remain entangled [^4] This issue has, in fact, already manifested in the discussion of a single Mössbauer wavepacket (Sec. 2-A).. Thus, the mathematical structure of general relativity—through the replacement of a universal time $t$ by different proper times $\tau_{A},\tau_{B}$—undermines the very notion of globally defined, simultaneously measurable EPR observables, exposing a tension between the EPR state and dilation of time in a weak gravitational field. Moreover, the entanglement of the $\lvert\mathrm{EPR}\rangle$ state eliminates the uncertainty of the combined variables, $\widehat{X}_{-}$ and $\widehat{P}_{+}$. Specifically:
$\widehat{x}_{A}-\widehat{x}_{B},\,\widehat{p}_{A}+\widehat{p}_{B}]=[\widehat{x}_{A},\widehat{p}_{A}]+[\widehat{x}_{A},\widehat{p}_{B}]-[\widehat{x}_{B},\widehat{p}_{A}]-[\widehat{x}_{B},\widehat{p}_{B}]=0.$
This can be described with quantum mechanical expectation values as:
$\left\langle\left[\Delta\left(\widehat{x}_{A}-\widehat{x}_{B}\right)\right]^{2}\right\rangle+\left\langle\left[\Delta\left(\widehat{p}_{A}+\widehat{p}_{B}\right)\right]^{2}\right\rangle=0.$
The meaning of this, as EPR articulated in their paper, is that nonlocal correlations allow these joint observables to evade the usual Heisenberg uncertainty constraints that apply to individual subsystems. In other words, if entanglement makes it possible to assign definite values to these combined variables without violating the commutation relations, then—according to the EPR reasoning—those variables must correspond to elements of physical reality.
The entanglement of the EPR state $\lvert\mathrm{EPR}\rangle$ can be described using either continuous or discrete bases, but it is not normalizable and therefore represents an unphysical quantum state. In practice, one uses the two-mode squeezed vacuum, denoted $\lvert\mathrm{EPR}^{\ast}\rangle$, which approaches the idealized $\lvert\mathrm{EPR}\rangle$ state asymptotically, reproducing its correlations arbitrarily well in the limit [9].
III.1 Proposed Experiment
As gravitational redshift offers the stage on which quantum mechanics is pitted against general relativity we propose to use photonic entanglement as an experimental testbed of local realism per EPR. In quantum mechanics a pair of entangled photons is viewed as a single system which maintains its correlation as long as it is not subject to collapse due to measurement. Even if the system undergoes decoherence due to gravity the lifetime of the correlation between the photons of the pair can be estimated by the coherence length:
$L_{c}=\frac{\lambda}{\left(\delta\lambda/\lambda\right)},\qquad\tau_{c}\approx\frac{L_{c}}{c}.$
Should the entangled pair obey the uncertainty relation, then, within the time estimated above, the correlation between the members of the pair should stay unperturbed. Hence, by testing if entanglement withstand gravitational redshift, or within what time it breaks down, we should be able to test the effect of a gravitational field on a quantum system. As depicted schematically in figure 2, Alice would be observing at ground level the interference pattern that arises between the laser beams A and A’, both with the same wavelengths.
While A’ is entangled with A”, A is entangled with B. Beam B undergoes redshift, while all the A beams do not. Bob’s measurement at a height $z$ (far enough from ground level) establishes that beam B is redshifted. The difference between the energy of Bob’s beam to that of Alice’s is, $\Delta E=\frac{h\nu}{c^{2}}\Delta\Phi_{G}$, where, $\Phi_{G}$, is the gravitational potential.
Figure 2: A general scheme of a proposed Mach–Zehnder interferometer for the experimental setup.
Alice’s measurement of beam A” serves as an indirect verification of the wavelength of beam A’, confirming its properties as initially generated. If beam B, which undergoes gravitational redshift, influences the wavelength of beam A due to entanglement, then the interference pattern formed between beams A and A’ will be modified. This would imply a nonlocal action at a distance—manifested as an observed change of the interference pattern. Thus, within a short time frame, Alice should be able to detect changes in the interference pattern between A and A’, induced by the gravitational redshift experienced by the entangled partner beam B.
The experiment could exhibit the following possible outcomes:
  1. 1.
    The interference pattern changes only after Bob performs his redshift measurement.
  2. 2.
    Bob doesn’t measure redshift on his beam but the interference pattern changes on Alice’s end within the above estimated lifetime.
  3. 3.
    The interreference pattern changes continuously from the onset of the experiment reaching a maximum around the time that it takes the beam to reach Bob’s location at $z$, after $z/c$ sec.
  4. 4.
    The interference pattern doesn’t change whether Bob performs his measurement or not.
While the first three outcomes indicate that entanglement influences spacetime and may therefore require incorporation into general relativity, the fourth outcome instead suggests that quantum action is fundamentally local and that GR does not require quantum degrees of freedom at all.
In addition, while Alice monitors the interference pattern on her end, spacetime may effectively function as a kind of “mindless Wigner’s friend” acting on beam B. That is, as beam B being redshifted while traveling upwards, entanglement may lead to a gradual degradation of the initially prepared interference pattern—or else the pattern may remain unaffected.
By contrast, when Bob measures the redshifted wavelength on his apparatus, he effectively becomes Alice’s “mindful Wigner’s friend,” as his conscious observation directly interacts with the quantum system [10]. If Bob’s measurement induces an abrupt change in the interference pattern, then his observation constitutes an active, conscious intervention in the entangled state.
Finally, it should also be noted that the laser beams can be polarized, which makes it possible to perform Bell-inequality measurements concurrently with the continuous-variable entanglement at both Alice’s and Bob’s stations. The results of these measurements would make it possible to compare the nature of these two completely different forms of entanglement.
Summary
We highlight a fundamental tension between the Heisenberg uncertainty principle and gravitational redshift, giving rise to the following paradox: how can gravitational redshift occur when a photon’s conjugate variables—momentum and position—cannot be simultaneously well defined in a manner compatible with local realism as articulated by EPR? To examine this issue, we propose a thought experiment employing continuous-variable quantum entanglement—an approach that has not yet been experimentally realized—in the presence of a weak gravitational field. The outcome of such an experiment may extend, and potentially empirically establish, the known incompatibility between quantum mechanics and general relativity into the weak-gravity regime.
By polarizing the beams, it becomes possible to perform Bell-inequality measurements concurrently with the continuous-variable entanglement at both ends of the setup. The results of these measurements would make it possible to compare the nature of these two fundamentally different forms of entanglement.
Acknowledgements
Asher Klatchko gratefully acknowledges David Griffiths, Emeritus Professor of Physics at Reed College, Robert Hamilton of George Fox University, Sandu Popescu of Bristol University and Daniel Rubén Zerzión of DIPC San Sebastián for insightful discussions and valuable comments.